OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 8x - y = 31 5x + 3y = 23

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

-y = -8x + 31 y = 8x -31 5x + 3(8x - 31) = 23 5x + 24x - 93 = 23 29x = 116 x = 4 8(4) - 31 = y y = 1

OpenStudy (anonymous):

y = 8x - 31 therefore 5x+24x-93 = 23 29x = 116 x=4 therefore y=1 (4,1)